OpenStudy (waheguru):

Can someone pleas ehelp me with this question, i am confused. Its solving porportions and i am doing 8:4:x = 2: y:6 how do i solve this?

OpenStudy (anonymous):

It turns out into a system of equations.

OpenStudy (anonymous):

\[ 8:4:x=c(2:y:6) \\ 8 = 2c \\ 4 = cy\\ x = 6c \\ \]

OpenStudy (anonymous):

First solve for c, then solve for x and y.

OpenStudy (waheguru):

where does the c come from?

OpenStudy (skullpatrol):

The alphabet.

OpenStudy (anonymous):

c is there to represent the fact that two proportional things are going to be equal if you multiply one of them by c.

OpenStudy (skullpatrol):

You need to find a c such that the above equation is true.

OpenStudy (waheguru):

I know a/b = c/d and a*d=b*c but this euqstion had 3 terms which is confusing me

OpenStudy (skullpatrol):

Have you solved for c?

OpenStudy (waheguru):

umm.

OpenStudy (waheguru):

im not sure how

OpenStudy (anonymous):

\[ 8=2c \]

OpenStudy (anonymous):

\[ \frac{8}{2} = \frac{4}{y} \\ \frac{8}{2} = \frac{x}{6} \]

OpenStudy (waheguru):

but how does 8 = 2c?

OpenStudy (skullpatrol):

$$8=2c$$ $$\frac{1}{2}*8=\frac{1}{2}*2c$$ $$4=(\frac{1}{2}*2)c$$ $$4=1c$$ $$4=c$$

OpenStudy (skullpatrol):

$$ 8:4:x = 2: y:6$$ $$8:4:x = c(2: y:6)$$

OpenStudy (skullpatrol):